The size of a television screen is given as the length of the screens diagonal. If the screens were flat, then the area of a square 21-inch screen would be how many square inches greater than the area of a square 19-inch screen???
Answer by edjones(7311)

Area of 21":
2s^2=21^2 Pythagoras
s^2=441/2
s=sqrt(220.5)
A=220.5
.
Area of 19":
2s^2=19^2
s^2=361/2
s=sqrt(180.5)
A=180.5
.
220.5-180.5=40 sq. in. difference
.
Ed